A train leaves a station and travels north at 60km/hr. Two hours later, a second train leaves on a parallel track and travels north at 100km/hr. How far from the station will they meet?
Answer by richwmiller(17219)

They will never meet. Parallel lines(tracks) never meet. Good thing too. We don;t need nor want a train accident!
When will they be the same distance from the original station?
Three hours after the second train leaves the station.
in the first two hours the first train traveled 2*60 =120 km
the second train gains on the first train at 100-60=40kph
120/40=3 hours